Annie runs a fitness center. On December 31, 2014, she bought an existing business with
exercise equipment and a building worth $300000. During 2015, business improved and
she bought some new equipment for $50000. At the end of 2015, her equipment and
buildings were worth $325000. Calculate Annie’s gross investment, depreciation, and net
investment during 2015.
1
Expert's answer
2020-03-26T10:11:48-0400

depreciationfor 2015"=(300000+50000)-325000=" $"25000"

gross investment for 2015 "=" $"50000"

net investment"=50000-25000="$ "25000"